OpenStudy (anonymous):

The perimeter of a rectangle is 90 cm. Its width is half its length. Which equation could be solved to find the dimensions of the rectangle?

OpenStudy (anonymous):

A. L + 1/2L = 90 B. 2L + 2 (1/2L) = 90 C. 2L + 2(L – 2) = 90 D. 2L + 2(L + 2) = 90

OpenStudy (anonymous):

I need help please I think its b but not too sure

OpenStudy (texaschic101):

P = 2L + 2W W = 1/2L P = 90 now just sub 90 = 2L + 2(1/2L) you are correct...it is B :) Nice job

OpenStudy (anonymous):

thank you